Systems Staff Meeting Notes 3/15/99

Systems Staff Meeting Notes
March 15, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Bobbie Meng, Everett Phillips, Suzy Shaw, Ying Tang
Announcements
  1. Everett placed 3rd in his age group in the Great Gainesville Road Race this weekend.
  2. Tickets will go on sale for WCW Monday Nitro Wrestling at the O'Connell Center (April 19).
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. We are in the process of unloading 20 cattle - 10 have already gone out, 5 more are scheduled for this afternoon. The next 20 that have been ordered will be stored here in Systems. Staff will be picking them up from this area. There is no distribution list yet for the next herd of 20.
  2. We've received 12 barcode readers, but have no distribution list yet.
  3. An order for yet another 20 Cows is in the works.
Liaisons
  1. Meeting this week is canceled (along with MMM) due to the Endeavor Demonstration at the Sheraton Hotel. Endeavor is the leading candidate in the OPAC race. The Directors are considering a new program to replace what is now NOTIS. Current quote on the project is $10 million.
Networking
  1. National DNS problems - at the level where they perform the registration of domain names - they were running an upgrade and thousands of names fell out of the registry. People took the opportunity to jump in & register, now they'll have to clean out the queue, which will result in disrupted communications.
  2. Preservation - the binding unit has been unable to send out email to a specific address. This could due to the DNS problems. There have also been general complaints of mail not being received from various sources. Expect a ripple effect throughout the week.
  3. Unix SP node problem at NERDC - is effecting Gatorlink and is causing access problems with WebLuis. Rich says that if the problem hasn't cleared up by the end of the day, he will temporarily replace the links on the homepage to work around the problem.
  4. Wiring estimate received - excluding the drops in this building and the network hardware for switches - comes in at about 10 grand.
Software Projects Status
  1. ATC Standalone - on hold pending Resource Services project that should be completed by this week - Marc record uploader. New specs are mushrooming, need to nip that in the bud.
  2. LASSIE - not yet ready for cloning. Documentation has been created & mailed out, Brian awaits feedback.
  3. InterDev - working on administrative aspects and have run into some roadblocks.
  4. Brian has been trying to set up Alena Aissing's student's database to access through Web interface. In SQL, it would require administrative access - this is not possible. Will try setting up single user access on SQL server.
  5. Ying is diligently working on the internal migration of the IR database.
  6. Mellon - awaiting feedback of beta testing from Harvard.
  7. We've received a query from the Health Center concerning the ATC software project.
DP Coordinator
  1. Compustat yearly tapes - documentation does not match the information on the tapes. Bill is trying to call Andy Halula at Compustat to have him point us to someone who has been able to read the tapes in an IBM format. We have been getting massive errors on both the ASCII & EBCDIC tapes.
Calendar
  1. Bill will be out most of the day on Thursday, March 18 - at the Endeavor demo.
  2. Brian will not be out the next two Thursdays.
  3. March 17 - St. Patrick's Day!
  4. Friday, March 19 - Bill & Suzy will both be out.
  5. Winston will be out March 29 - April 2.
  6. Everett will be out April 2 and April 5.
  7. Rich will be out April 6-9.


Last updated February 18, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu